(u^8z^7 ) / (z^3u^ -2 )
I have that this can't be simplied anymore. I think I'm correct just want to make sure.
36^1/2/15^2
I'm really lost on this one I want to go ahead a ^2 15 and get 225, but I come put with a decimal as the answer.

Answer by checkley77(12844)

(u^8z^7 ) / (z^3u^ -2 )
U^(8+2)Z^(7-3)=U^10Z^4 ANSWER.
--------------------------------------
36^1/2/15^2
SQRT36/225
6/225=2/75 ANSWER.